# Streamlined WZ method proofs of Van Hamme supercongruences

### v2

@@ 1. Introduction
Many authors have extended the original Van Hamme supercongruences to hold modulo higher powers of $p$, see for example Swisher [24]. Liu extended H.2 for primes $p\equiv 3\pmod{4}$ [14]. Recently, Guo and Wang used the WZ method to prove E.2 and F.2 [7]. The proof given here is different, and in some ways simpler because the right hand side of the WZ pair telescopes to zero. That is, the proofs here are similar to Zudilin’s proof of B.2 [26], rather than to Osburn and Zudilin’s proof of K.2 [21] which telescopes the left hand side. In addition, the present proofs of B.2-H.2 have the same shape. However, Guo and Wang’s proof [7] is more general and applies to more cases. In another recent development, Jana and Karmakar prove conjectures of Guo related to B.2 and C.2 using the WZ method and a parametrized WZ pair [11]. In their work, they state that despite this generalization the WZ method still requires an inspired guess. In contrast, the approach presented here introduces the notion of a WZ device with the intent of mechanically recovering a suitable WZ pair from the output of the WZ algorithm. This approach is somewhat similar to that of Guillera for infinite series [3].
@@
The remainder of the paper is organized as follows. Section 2 summarizes background material and auxiliary results. This section also shows the Van Hamme supercongruence I.2 can be proved as a special case of the WZ method. Section 3 develops a procedure to find candidates for the clever factors needed for the WZ method to succeed, and exemplifies it to prove H.2. In Section 4, the preceding results together with an important remark are used to prove the supercongruences B.2, C.2, E.2, F.2, and G.2 with a uniform proof structure. In particular, this method yields the supercongruence G.2 modulo $p^{4}$ as shown in Swisher [24], and H.2 modulo $p^{3}$ when $p\equiv 3\pmod{4}$ as shown in Liu [14].
To develop these ideas further, Section 5 introduces the notion of a WZ device. This section also proves D.2, completing the proof of Theorem 1.2. Each of the proofs given in Sections 3–5 utilizes a WZ device. A list of WZ devices used here is provided at the end of the paper in Table 2. The proofs for C.2 and D.2 suggest they hold modulo higher powers of $p$, as noted by Swisher [24] and Long and Ramakrishna [17], respectively. An argument in terms of the present approach remains elusive.
@@ 2.2. Long-Ramakrishna approximation for $\Gamma_{p}$
While working with congruences, the following sections implicitly rely on [17, Proposition 13], which states that for primes $p\geq 5$, integers $0\leq k<p$, and $a\in\mathbb{Q}$,
(14) $\operatorname{val}_{p}(G_{k}(a)/k!)=0.$
The next theorem is a key observation about congruences involving $\Gamma_{p}$, and combines the results in Long and Ramakrishna [17, Thm. 14], and Swisher [24, Eqn. (15)].
@@ 2.3. Auxiliary results
The denominator above does not have factors of $p$ because of (14). Thus, these two factors cancel modulo $p^{3}$, as desired. ∎
@@ 5. WZ devices
Multiplying these factors together forms a fraction. Compute this fraction modulo $p^{3}$, then apply (12) to make the evaluations of $G_{1}$ match to find both numerator and denominator are

### v3

@@ 1. Introduction
Many authors have extended the original Van Hamme supercongruences to hold modulo higher powers of $p$, see for example Swisher [24]. Liu extended (H.2) for primes $p\equiv 3\pmod{4}$ [14]. Recently, Guo and Wang used the WZ method to prove (E.2) and (F.2) [4]. The proof given here is different, and in some ways simpler because the right hand side of the WZ pair telescopes to zero. That is, the proofs here are similar to Zudilin’s proof of (B.2) [26], rather than to Osburn and Zudilin’s proof of (K.2) [21] which telescopes the left hand side. In addition, the present proofs of (B.2)-(H.2) have the same shape. However, Guo and Wang’s proof [4] is more general and applies to more cases. In another recent development, Jana and Karmakar prove conjectures of Guo related to (B.2) and (C.2) using the WZ method and a parametrized WZ pair [11]. In their work, they state that despite this generalization the WZ method still requires an inspired guess. In contrast, the approach presented here introduces the notion of a WZ device with the intent of mechanically recovering a suitable WZ pair from the output of the WZ algorithm. This approach is somewhat similar to that of Guillera for infinite series [3].
@@
The remainder of the paper is organized as follows. Section 2 summarizes background material and auxiliary results. This section also shows the Van Hamme supercongruence (I.2) can be proved as a special case of the WZ method. Section 3 develops a procedure to find candidates for the clever factors needed for the WZ method to succeed, and exemplifies it to prove (H.2). In Section 4, the preceding results together with an important remark are used to prove the supercongruences (B.2), (C.2), (E.2), (F.2), and (G.2) with a uniform proof structure. In particular, this method yields the supercongruence (G.2) modulo $p^{4}$ as shown in Swisher [24], and (H.2) modulo $p^{3}$ when $p\equiv 3\pmod{4}$ as shown in Liu [14].
To develop these ideas further, Section 5 introduces the notion of a WZ device. This section also proves (D.2), completing the proof of Theorem 1.2. Each of the proofs given in Sections 3–5 utilizes a WZ device. A list of WZ devices used here is provided at the end of the paper in Table 2. The proofs for (C.2) and (D.2) suggest they hold modulo higher powers of $p$, as noted by Swisher [24] and Long and Ramakrishna [16], respectively. An argument in terms of the present approach remains elusive.
@@ 2.2. Long-Ramakrishna approximation for $\Gamma_{p}$
The next theorem is a key observation about congruences involving $\Gamma_{p}$, and combines the results in Long and Ramakrishna [16, Thm. 14], and Swisher [24, Eqn. (15)].
@@ 2.3. Auxiliary results
The denominator above does not have factors of $p$. Thus, these two factors cancel modulo $p^{3}$, as desired. ∎
@@ 5. WZ devices
Multiplying these factors together forms a quotient. Compute this quotient modulo $p^{3}$, then apply (12) to make the evaluations of $G_{1}$ match to find both numerator and denominator are